# EHRStruct: A Comprehensive Benchmark Framework for Evaluating Large Language Models on Structured Electronic Health Record Tasks

### v1

@@ 3.2 Task Taxonomy Construction
As summarized in Table 1, we organize the benchmark tasks along three axes: clinical scenario (Data-Driven vs. Knowledge-Driven), cognitive level (Understanding vs. Reasoning), and task category (six functional types). This taxonomy captures both the practical intent and the reasoning complexity of each task, supporting a comprehensive and interpretable evaluation framework. Detailed descriptions are provided in Appendix A.2, and task instructions appear in Appendix A.3. Such a structured organization also enables fine-grained comparison of model performance across clinically and cognitively diverse settings.
@@ 3.3 Task-Specific Samples Extraction
GPT3.5turbo (openai2023) Commercial
GPT4.1 (openai2025) Commercial
@@
DeepSeekV2.5 (liu2024deepseeka) 236B
DeepSeekV3 (liu2024deepseekb) 685B
Qwen7B (team2024qwen2) 7B
Qwen14B (team2024qwen2) 14B
Qwen32B (team2024qwen2) 32B
Qwen72B (team2024qwen2) 72B
@@
Meditron7B (chen2023meditron) 7B
MedAlpaca13B (yuan2024continued) 13B
@@ 3.6 EHRMaster Overview
Here, we briefly outline our proposed EHRMaster framework. A detailed description is provided in Section B. EHRMaster operates in three key stages:
@@ D-U1 D-U2 D-R1 D-R2 D-R3 D-R4 D-R5 K-U1 K-R1 K-R2 K-R3
DeepSeekV2.5 72 41 18 51 14 44 52 51
DeepSeekV3 72 41 8 37 12 72 90 52.8
Qwen7B 1 7 4 24 1
Qwen14B 4 30 19 17 11 16 4
Qwen32B 25 25 24 26 15 47 10 58.3 51
Qwen72B 15 6 27 48 20 41 29 52.2
@@ 4.4 Finetuning Analysis
To analyze the fine-tuning performance of large language models on structured EHR tasks, we construct additional training samples specifically for fine-tuning, which are completely separate from the evaluation samples used in benchmarking. Each fine-tuning dataset contains 30 task-specific question–answer table pairs following the same instruction format as in the evaluation, ensuring consistency of task structure while preventing data leakage. We adopt two fine-tuning strategies: single-task fine-tuning trains each task separately, while multi-task fine-tuning combines all tasks into one set. All experiments are conducted on the Qwen7B model using LoRA-based parameter-efficient fine-tuning with a 10
@@ C.4 Full Benchmark Results on Synthea and eICU
In contrast, most medical-domain LLMs fail to produce valid outputs in the Knowledge-Driven setting. On the Synthea dataset, models such as Huatuo, HEAL, Meditron7B, and PMCLLaMA13B yield no valid output on all four Knowledge-Driven tasks. Their performance in Data-Driven tasks is also low: for instance, MedAlpaca13B scores no higher than 11 on any task, and Med4270B reaches only 27 on D-R4. Similar patterns hold on the eICU dataset, where most medical models again fail to output any valid AUC and achieve accuracy below 25% across nearly all Data-Driven tasks. None of the medical-domain models ranks in the top-3 for any task in either dataset.
@@
While this reasoning-related gap is consistent across datasets, the absolute performance also heavily depends on the dataset. For instance, DeepSeekV3 scores 59 on D-R5 in eICU, compared to 90 in Synthea, showing large variance even under the same task design. These results suggest that real-world data introduces additional modeling challenges beyond reasoning complexity alone.
@@
Across both the Synthea and eICU datasets, nearly all medical-domain LLMs fail to produce valid outputs on Knowledge-Driven tasks. Specifically, models such as Huatuo, HEAL, Meditron7B, PMCLLaMA13B, and MedAlpaca13B output no valid AUC on any of the four Knowledge-Driven tasks in either dataset. Even the relatively stronger medical models, such as ClinicalCamel70B and Med4270B, produce only one or two valid outputs across all Knowledge-Driven tasks, and their AUC scores remain below 50. In contrast, general-purpose LLMs like GPT-4.1 and Gemini-2.5 not only produce valid predictions on all Knowledge-Driven tasks but also achieve AUCs consistently above 50, with top scores reaching 62.1 on K-R3 in eICU.
@@ D-U1 D-U2 D-R1 D-R2 D-R3 D-R4 D-R5 K-U1 K-R1 K-R2 K-R3
GPT3.5turbo plain text conversion 6 15 14 18 7 7 24 58.1 55.4 52.9
GPT3.5turbo special character separation 8 14 20 25 18 3 7 51 59.3 52.2 60.2
GPT3.5turbo graph-structured representation 7 28 10 19 12 3 8 53.1 53.1
GPT3.5turbo natural language description 5 16 15 26 14 11 59 50.9 57.7
GPT4.1 plain text conversion 79 51 52 56 48 70 84 55 55.6 53.2 51
GPT4.1 special character separation 75 50 43 50 43 65 70 53.3 54.5 55.6
GPT4.1 graph-structured representation 58 54 48 60 47 66 80 54 50.7
GPT4.1 natural language description 78 51 48 57 45 68 81 55 54.9
@@
DeepSeekV2.5 plain text conversion 72 41 18 51 14 44 52 51
DeepSeekV2.5 special character separation 72 37 13 42 15 47 56 52.8
DeepSeekV2.5 graph-structured representation 26 19 14 41 13 28 42 52
DeepSeekV2.5 natural language description 65 39 13 49 16 74 73 54.1
DeepSeekV3 plain text conversion 72 41 8 37 12 72 90 52.8 56.4
DeepSeekV3 special character separation 66 40 9 42 10 69 84 55.6
DeepSeekV3 graph-structured representation 72 37 13 47 14 76 89 55.6
DeepSeekV3 natural language description 65 39 11 38 12 88 96 52.8 51.0
Qwen7B plain text conversion 1 7 4 24 1
Qwen7B special character separation 1 8 4 20 3
Qwen7B graph-structured representation 1 12 4 19 2
Qwen7B natural language description 2 9 8 16 4 12 13
Qwen14B plain text conversion 4 30 19 17 11 16 4
Qwen14B special character separation 5 29 17 14 11 16 6 52
Qwen14B graph-structured representation 5 28 14 21 9 13 2
Qwen14B natural language description 2 22 18 13 8 44 35
Qwen32B plain text conversion 25 25 24 26 15 47 10 58.3 51
Qwen32B special character separation 26 23 26 32 16 51 18 55.6 51
Qwen32B graph-structured representation 38 25 20 28 14 58 14 58.3 51
Qwen32B natural language description 32 22 22 35 16 72 45 55.6 51
Qwen72B plain text conversion 15 6 27 48 20 41 29 52.2
Qwen72B special character separation 18 6 23 42 16 46 34
Qwen72B graph-structured representation 16 7 29 39 25 31 17
Qwen72B natural language description 21 9 27 31 18 74 64
@@
DeepSeekV2.5 72 41 18 47 14 44 52 53.2
DeepSeekV3 72 41 8 61 12 72 59 52.8
Qwen7B 1 7 4 24 1
Qwen14B 4 30 19 17 11 16 4
Qwen32B 25 25 24 26 15 47 10 58.3 51
Qwen72B 15 6 27 48 20 41 29 52.2

### v3

@@ 3.2 Task Taxonomy Construction
As summarized in Table 1, we organize the benchmark tasks along three axes: clinical scenario (Data-Driven vs. Knowledge-Driven), cognitive level (Understanding vs. Reasoning), and task category (six functional types). This taxonomy captures both the practical intent and the reasoning complexity of each task, supporting a comprehensive and interpretable evaluation framework. Detailed descriptions are provided in Appendix Section A.2, and task instructions appear in Appendix Section A.3. Such a structured organization also enables fine-grained comparison of model performance across clinically and cognitively diverse settings.
@@ 3.3 Task-Specific Samples Extraction
GPT-3.5 Turbo (openai2023) Commercial
GPT-4.1 (openai2025) Commercial
@@
DeepSeek-V2.5 (liu2024deepseeka) 236B
DeepSeek-V3 (liu2024deepseekb) 685B
Qwen-7B (team2024qwen2) 7B
Qwen-14B (team2024qwen2) 14B
Qwen-32B (team2024qwen2) 32B
Qwen-72B (team2024qwen2) 72B
@@
Meditron-7B (chen2023meditron) 7B
MedAlpaca-13B (yuan2024continued) 13B
@@ 3.6 EHRMaster Overview
Here, we briefly outline our proposed EHRMaster framework. A detailed description is provided in Appendix Section B. EHRMaster operates in three key stages:
@@ D-U1 D-U2 D-R1 D-R2 D-R3 D-R4 D-R5 K-U1 K-R1 K-R2 K-R3
DeepSeek-V2.5 72 41 18 51 14 44 52 51
DeepSeek-V3 72 41 8 37 12 72 90 52.8
Qwen-7B 1 7 4 24 1
Qwen-14B 4 30 19 17 11 16 4
Qwen-32B 25 25 24 26 15 47 10 58.3 51
Qwen-72B 15 6 27 48 20 41 29 52.2
@@ 4.4 Finetuning Analysis
To analyze the fine-tuning performance of large language models on structured EHR tasks, we construct additional training samples specifically for fine-tuning, which are completely separate from the evaluation samples used in benchmarking. Each fine-tuning dataset contains 30 task-specific question–answer table pairs following the same instruction format as in the evaluation, ensuring consistency of task structure while preventing data leakage. We adopt two fine-tuning strategies: single-task fine-tuning trains each task separately, while multi-task fine-tuning combines all tasks into one set. All experiments are conducted on the Qwen-7B model using LoRA-based parameter-efficient fine-tuning with a 10% validation split, a learning rate of 0.0001, three training epochs, a batch size of 8, and LoRA hyperparameters set to rank 8, alpha 32, and dropout 0.05.
@@ 5 Conclusion
In future work, we plan to extend EHRStruct to evaluate iterative and adaptive treatment-planning processes that incorporate real-time patient responses such as drug reactions and physiological indicators, better reflecting the dynamic nature of clinical decision-making.
@@ C.4 Full Benchmark Results on Synthea and eICU
In contrast, most medical-domain LLMs fail to produce valid outputs in the Knowledge-Driven setting. On the Synthea dataset, models such as Huatuo, HEAL, Meditron-7B, and PMC_LLaMA_13B yield no valid output on all four Knowledge-Driven tasks. Their performance in Data-Driven tasks is also low: for instance, MedAlpaca-13B scores no higher than 11 on any task, and Med42-70B reaches only 27 on D-R4. Similar patterns hold on the eICU dataset, where most medical models again fail to output any valid AUC and achieve accuracy below 25% across nearly all Data-Driven tasks. None of the medical-domain models ranks in the top-3 for any task in either dataset.
@@
While this reasoning-related gap is consistent across datasets, the absolute performance also heavily depends on the dataset. For instance, DeepSeek-V3 scores 59 on D-R5 in eICU, compared to 90 in Synthea, showing large variance even under the same task design. These results suggest that real-world data introduces additional modeling challenges beyond reasoning complexity alone.
@@
Across both the Synthea and eICU datasets, nearly all medical-domain LLMs fail to produce valid outputs on Knowledge-Driven tasks. Specifically, models such as Huatuo, HEAL, Meditron-7B, PMC_LLaMA_13B, and MedAlpaca-13B output no valid AUC on any of the four Knowledge-Driven tasks in either dataset. Even the relatively stronger medical models, such as ClinicalCamel70B and Med42-70B, produce only one or two valid outputs across all Knowledge-Driven tasks, and their AUC scores remain below 50. In contrast, general-purpose LLMs like GPT-4.1 and Gemini-2.5 not only produce valid predictions on all Knowledge-Driven tasks but also achieve AUCs consistently above 50, with top scores reaching 62.1 on K-R3 in eICU.
@@ D-U1 D-U2 D-R1 D-R2 D-R3 D-R4 D-R5 K-U1 K-R1 K-R2 K-R3
GPT-3.5 Turbo plain text conversion 6 15 14 18 7 7 24 58.1 55.4 52.9
GPT-3.5 Turbo special character separation 8 14 20 25 18 3 7 51 59.3 52.2 60.2
GPT-3.5 Turbo graph-structured representation 7 28 10 19 12 3 8 53.1 53.1
GPT-3.5 Turbo natural language description 5 16 15 26 14 11 59 50.9 57.7
GPT-4.1 plain text conversion 79 51 52 56 48 70 84 55 55.6 53.2 51
GPT-4.1 special character separation 75 50 43 50 43 65 70 53.3 54.5 55.6
GPT-4.1 graph-structured representation 58 54 48 60 47 66 80 54 50.7
GPT-4.1 natural language description 78 51 48 57 45 68 81 55 54.9
@@
DeepSeek-V2.5 plain text conversion 72 41 18 51 14 44 52 51
DeepSeek-V2.5 special character separation 72 37 13 42 15 47 56 52.8
DeepSeek-V2.5 graph-structured representation 26 19 14 41 13 28 42 52
DeepSeek-V2.5 natural language description 65 39 13 49 16 74 73 54.1
DeepSeek-V3 plain text conversion 72 41 8 37 12 72 90 52.8 56.4
DeepSeek-V3 special character separation 66 40 9 42 10 69 84 55.6
DeepSeek-V3 graph-structured representation 72 37 13 47 14 76 89 55.6
DeepSeek-V3 natural language description 65 39 11 38 12 88 96 52.8 51.0
Qwen-7B plain text conversion 1 7 4 24 1
Qwen-7B special character separation 1 8 4 20 3
Qwen-7B graph-structured representation 1 12 4 19 2
Qwen-7B natural language description 2 9 8 16 4 12 13
Qwen-14B plain text conversion 4 30 19 17 11 16 4
Qwen-14B special character separation 5 29 17 14 11 16 6 52
Qwen-14B graph-structured representation 5 28 14 21 9 13 2
Qwen-14B natural language description 2 22 18 13 8 44 35
Qwen-32B plain text conversion 25 25 24 26 15 47 10 58.3 51
Qwen-32B special character separation 26 23 26 32 16 51 18 55.6 51
Qwen-32B graph-structured representation 38 25 20 28 14 58 14 58.3 51
Qwen-32B natural language description 32 22 22 35 16 72 45 55.6 51
Qwen-72B plain text conversion 15 6 27 48 20 41 29 52.2
Qwen-72B special character separation 18 6 23 42 16 46 34
Qwen-72B graph-structured representation 16 7 29 39 25 31 17
Qwen-72B natural language description 21 9 27 31 18 74 64
@@
DeepSeek-V2.5 72 41 18 47 14 44 52 53.2
DeepSeek-V3 72 41 8 61 12 72 59 52.8
Qwen-7B 1 7 4 24 1
Qwen-14B 4 30 19 17 11 16 4
Qwen-32B 25 25 24 26 15 47 10 58.3 51
Qwen-72B 15 6 27 48 20 41 29 52.2